# DePT3R: Joint Dense Point Tracking and 3D Reconstruction of Dynamic Scenes in a Single Forward Pass

### v1

@@ Abstract
Current methods for dense 3D point tracking in dynamic scenes typically rely on pairwise processing, require known camera poses, or assume a temporal ordering to input frames, constraining their flexibility and applicability. Additionally, recent advances have successfully enabled efficient 3D reconstruction from large-scale, unposed image collections, underscoring opportunities for unified approaches to dynamic scene understanding. Motivated by this, we propose DePT3R, a novel framework that simultaneously performs dense point tracking and 3D reconstruction of dynamic scenes from multiple images in a single forward pass. This multi-task learning is achieved by extracting deep spatio-temporal features with a powerful backbone and regressing pixel-wise maps with dense prediction heads. Crucially, DePT3R operates without requiring camera poses, substantially enhancing its adaptability and efficiency—especially important in dynamic environments with rapid changes. We validate DePT3R on several challenging benchmarks involving dynamic scenes, demonstrating strong performance and significant improvements in memory efficiency over existing state-of-the-art methods. Data and codes are available via the open repository: https://github.com/StructuresComp/DePT3R
@@ 1 Introduction
Method
Dynamic
Scene
Unposed
Dense Point
Tracking
Temporal
Attention
Point Tracking 3D Reconstruction
@@ 2.1 Tabula Rasa 3D Reconstruction
Tabula rasa methods aim to reconstruct 3D scenes solely from raw input observations without reliance on prior scene knowledge. Early techniques, such as Structure from Motion (SfM), relied heavily on matching handcrafted features across multiple source images to estimate camera poses and reconstruct scene structure [10]. The advent of Neural Radiance Fields (NeRF) [20] initiated a paradigm shift by formulating reconstruction as an optimization-based novel-view synthesis problem, where scenes are represented as continuous neural fields. InstantNGP [21] accelerated this optimization by coupling a compact neural network with a multi-resolution hash table. Further improvements in rendering speed were achieved by 3D Gaussian Splatting [16], which explicitly represents scenes as collections of anisotropic Gaussian functions, efficiently rasterized for real-time rendering. Subsequently, Mip-Splatting [42] reduced rendering aliasing by constraining Gaussian sizes. Despite impressive results in view synthesis, these approaches necessitate per-scene optimization and cannot leverage learned data priors.
@@ 2.2 Learning-based 3D Reconstruction
One notable breakthrough was DUSt3R [36], which enabled 3D reconstruction from unposed image pairs using asymmetric transformers linked through cross-attention mechanisms. The predictions from individual image pairs were subsequently globally optimized in a standard coordinate system. Subsequent developments included symmetric architectures and global attention layers, which significantly improved memory efficiency for processing larger image sets [40].
@@
Estimation of 3D motion of points has also advanced significantly. RAFT-3D [29] employed rigid motion embeddings to estimate scene flow between RGBD pairs. Omnimotion [34] unified 2D and 3D tracking with a quasi-3D canonical representation but still required per-scene optimization. Introduced by Xiao et al., SpatialTracker [38] enables feed-forward tracking by creating a triplane scene encoding with depth estimation. The core mechanism is a transformer that iteratively refines query point trajectories, producing updated paths using the prior trajectory, the query point’s features, and the features of nearby points. DELTA [22] efficiently computed dense 3D trajectories, while TAPIP3D [43] leveraged depth to elevate image features into a global coordinate system for enhanced performance. Building on the DUSt3R architecture, St4RTrack [7] adapted it for 3D point tracking by incorporating an additional prediction head and chaining motion predictions for longer trajectories. However, its pairwise processing limits the exploitation of temporal attention, thereby restricting its ability to handle significant camera motion effectively. POMATO [45] addressed these limitations by introducing a temporal attention module, eliminating scale normalization, and enhancing inter-frame interactions. Nonetheless, its fundamentally pairwise processing strategy continued to restrict fully global interactions, indicating room for further advancement. Stereo4D  [12] attached a temporal motion head to the DUSt3R architecture used to estimate the position of scene points at any query time between the time of the two input frames and introduced a large dataset of internet videos with noisy point trajectories to train the motion head.
@@ 3.1 Preliminary
Conversely, in a dynamic scene, most pointmaps cannot be described by simple geometric transforms alone, as they inherently encode scene motion. Note that by leveraging $\prescript{c}{}{\mathbf{X}}^{a}_{t_{1}}$ and $\prescript{c}{}{\mathbf{X}}^{a}_{t_{2}}$, we can effectively analyze the motion of the scene points visible in frame $a$ from time of frame $t_{1}$ to the time of frame $t_{2}$.
@@
In the VGGT method, each input image is first decomposed into image tokens via DINOv2  [23]. These tokens, along with a learnable camera token and four learnable register tokens per image, are then input to a global aggregator module. Since predictions are made in the initial frame’s coordinate system, the camera and register tokens for the first image differ from those for subsequent frames, enabling the network to discriminate among frames. The aggregator module consists of alternating local and global attention layers: local layers facilitate interactions among tokens within the same image, while global layers promote interactions across images.
@@ 3.2 Point Tracking and Dynamic Reconstruction
Our method takes a sequence of $N$ frames, $(I_{t})^{N}_{t=1}$, and a query time $q$, and performs joint reconstruction of the scene and 3D tracking of observable points. To this end, for each frame $I_{t}$, we predict two pixel-wise maps expressed in the coordinate system of the first frame $I_{1}:\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{t}$, capturing point positions at the observation time, and $\prescript{1}{}{\mathbf{\hat{M}}}^{t}_{q}=\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{q}-\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{t}$, capturing the motion of points from time of frame $t$ to the time of query frame $q$. By jointly estimating these maps, our method concurrently reconstructs the 3D structure and tracks point motion to the query frame. Notably, our tracking approach establishes direct correspondences between each frame and the query frame, rather than using pairwise, frame-to-frame tracking, which enables the efficient capture of overall motion and deformation over extended intervals.
@@ 4.1 Experimental Setups
where $\delta_{3D}=\{\mathrm{0.1m,0.3m,0.5m,1.0m}\}$. Since predictions are made in the coordinate system of the first frame’s camera, we align the estimated 3D positions with the ground truth by scaling them by the ratio of the median of the norms of the estimated and ground-truth positions. That is, we multiply the predicted coordinates by the $s$ factor:
@@
During training, sequences of 2 to 10 frames are randomly selected from all available sequences. For Kubric Movi-F, we use a stride of $1$; for all other datasets, we use a stride varying between $1$ and $4$. TartanAir and Virtual KITTI 2 are sampled with half the frequency of the different datasets.
We observe that training with the final tracking objective immediately, using model weights initialized with VGGT’s weights, degrades performance. Thus, we perform training in two phases. In the first phase, we train DePT3R using only the camera, depthmap, and pointmap losses. The intrinsic embedding is used in the first phase, but the query embedding is not. In the second phase, we integrate the query embedding and motion head, along with the tracking loss. To improve training stability and speed, we do not use a confidence-based loss in the second phase.
We employ common data augmentation strategies, such as color jitter, random aspect ratios, and random center crops  [36], applied uniformly over each image of the sequence. We use the Adam optimizer with a warmup phase, followed by cosine scheduling. Gradient checkpointing and bfloat16 precision are utilized to enhance efficiency and reduce GPU memory usage. In the first phase, we train on all datasets, using a learning rate of $5\times 10^{-5}$ for the intrinsic embedding and a learning rate of $5\times 10^{-6}$ for the other weights. In the second phase, we train only on PointOdyssey, DynamicReplica, and Kubric, using a learning rate of $1\times 10^{-5}$ for the motion head and the query embedding, and a learning rate of $1\times 10^{-6}$ for all other weights. Both training phases are done on a single NVIDIA RTX 6000 and collectively take approximately three days. We use PyTorch 2.3.1 as our machine learning framework. Experiments are performed on a machine equipped with an AMD Ryzen Threadripper PRO 7975WX 32-core CPU and Ubuntu 22.04 installed.
@@ 4.2 Experimental Results
3D Point Tracking. We first evaluate the 3D point tracking task across three datasets: the test sets of the synthetic datasets PointOdyssey and DynamicReplica, which feature significant camera and scene motion, and PanopticStudio, a real-world dataset with no camera motion. For evaluation, we followed the same setup of St4RTrack [7] and selected 64 consecutive frames from 50 randomly selected sequences in each dataset. To compare with previous work, we restrict our evaluation to the trajectories of the points from the first frame, even though our method can track all points. Following St4RTrack, we exclude sequences from PointOdyssey that contain fog or are generated in the Kubric style.
As shown in Table 2, our method significantly outperforms all baselines across all datasets. Remarkably, even though our model was trained only on sequences of at most 10 images, it achieves impressive performance on much longer sequences. Moreover, unlike other approaches that require a windowing strategy or global optimization to handle extended sequences due to memory or architecture constraints, our method can generate predictions for a large number of frames in a single forward pass.
3D Reconstruction. Following St4RTrack [7], we diverge from previous work, directly evaluating the accuracy of our estimated 3D reconstruction with the APD and EPE metrics. We use the TUM RGB-D Benchmark, a real-world dataset with significant camera motion, scene motion, and motion blur, as well as the test set of PointOdyssey. We report metrics on 50 randomly chosen sequences of 64 consecutive frames. For the TUM RGB-D Benchmark, we filter out points with depths between 0.1 and 5 meters, as the depth camera’s accuracy degrades at long ranges. On the TUM RGB-D Benchmark (Table 3), DePT3R achieves an APD of 92.22 and an EPE of 0.0968, competitive with VGGT and significantly better than the other baselines. On PointOdyssey, DePT3R achieves an EPE of 0.0406 and an APD of 98.01, noticeably better than all of the baselines. These results demonstrate that DePT3R achieves state-of-the-art 3D reconstruction accuracy on both synthetic and real-world data, including out-of-distribution scenarios, using synthetic training data.
@@
Figure 4: Qualitative ablation study on the Panoptic Studio dataset. The 3D trajectories of scene points are projected onto the last frame. Without the intrinsic embedding, the model can identify the direction of motion but struggles to accurately place it and gauge its magnitude, resulting in significant errors in point trajectory estimation.
@@
Figure 5: GPU memory usage comparison between SpatialTrackerV2, VGGT and our DePT3R method across varying numbers of query points. SpatialTrackerV2 and VGGT exhibit a rapid increase in GPU memory consumption, exhausting the 48 GB memory limit at just 40k and 22.5k query points, respectively. In contrast, DePT3R efficiently handles 268k query points, requiring only 12 GB of memory. All experiments were performed on an RTX A6000 GPU (48 GB). Zoom in for better visualization.
@@ 5 Conclusion and Future Works
Summary. In this paper, we introduce DePT3R, a simple, novel framework for simultaneously performing dense 3D reconstruction and point tracking of dynamic scenes from unposed monocular video. Despite training exclusively on synthetic datasets, our approach achieves strong performance on challenging, real-world benchmarks, highlighting the inherent synergy between reconstruction and point tracking tasks. Unlike prior methods, DePT3R leverages global temporal attention to enhance both accuracy and computational efficiency, enabling dense tracking of all visible points in a single, computationally efficient forward pass. Additionally, we effectively leverage intrinsic information to improve reconstruction and motion estimation significantly.
@@ Acknowledgements
This work was supported in part by the US Department of Agriculture (Grant numbers 2024-67021-42528 and 2022-67022-37021).

### v2

@@ Abstract
Current methods for dense 3D point tracking in dynamic scenes typically rely on pairwise processing, require known camera poses, or assume temporal ordering of input frames, thereby constraining their flexibility and applicability. Additionally, recent advances have successfully enabled efficient 3D reconstruction from large-scale, unposed image collections, underscoring opportunities for unified approaches to dynamic scene understanding. Motivated by this, we propose DePT3R, a novel framework that simultaneously performs dense point tracking and 3D reconstruction of dynamic scenes from multiple images in a single forward pass. This multi-task learning is achieved by extracting deep spatio-temporal features with a powerful backbone and regressing pixel-wise maps with dense prediction heads. Crucially, DePT3R operates without requiring camera poses, substantially enhancing its adaptability and efficiency—especially important in dynamic environments with rapid changes. We validate DePT3R on several challenging benchmarks involving dynamic scenes, demonstrating strong performance and significant improvements in memory efficiency over existing state-of-the-art methods. Data and codes are available via the open repository: https://github.com/StructuresComp/DePT3R
@@ 1 Introduction
Method Dynamic Scene Unposed Dense Point Tracking Temporal Attention Point Tracking 3D Reconstruction
@@ 2.1 Tabula Rasa 3D Reconstruction
Tabula rasa methods aim to reconstruct 3D scenes solely from raw input observations, without relying on prior scene knowledge. Early techniques, such as Structure from Motion (SfM), relied heavily on matching handcrafted features across multiple source images to estimate camera poses and reconstruct scene structure [10]. The advent of Neural Radiance Fields (NeRF) [20] initiated a paradigm shift by formulating reconstruction as an optimization-based novel-view synthesis problem, where scenes are represented as continuous neural fields. InstantNGP [21] accelerated this optimization by coupling a compact neural network with a multi-resolution hash table. Further improvements in rendering speed were achieved by 3D Gaussian Splatting [16], which explicitly represents scenes as collections of anisotropic Gaussian functions, efficiently rasterized for real-time rendering. Subsequently, Mip-Splatting [42] reduced rendering aliasing by constraining Gaussian sizes. Despite impressive results in view synthesis, these approaches necessitate per-scene optimization and cannot leverage learned data priors.
@@ 2.2 Learning-based 3D Reconstruction
One notable breakthrough was DUSt3R [36], which enabled 3D reconstruction from unposed image pairs using asymmetric transformers linked via cross-attention. The predictions from individual image pairs were subsequently globally optimized in a standard coordinate system. Subsequent developments included symmetric architectures and global attention layers, which significantly improved memory efficiency for processing larger image sets [40].
@@
Estimation of 3D point motion has also advanced significantly. RAFT-3D [29] employed rigid motion embeddings to estimate scene flow between RGBD pairs. Omnimotion [34] unified 2D and 3D tracking with a quasi-3D canonical representation but still required per-scene optimization. Introduced by Xiao et al. [38], SpatialTracker enables feed-forward tracking by creating a triplane scene encoding with depth estimation. The core mechanism is a transformer that iteratively refines query point trajectories, producing updated paths using the prior trajectory, the query point’s features, and the features of nearby points. DELTA [22] efficiently computed dense 3D trajectories, while TAPIP3D [43] leveraged depth to elevate image features into a global coordinate system for enhanced performance. Building on the DUSt3R architecture, St4RTrack [7] adapted it for 3D point tracking by incorporating an additional prediction head and chaining motion predictions for longer trajectories. However, its pairwise processing limits the exploitation of temporal attention, thereby restricting its ability to handle significant camera motion effectively. POMATO [45] addressed these limitations by introducing a temporal attention module, eliminating scale normalization, and enhancing inter-frame interactions. Nonetheless, its fundamentally pairwise processing strategy continued to restrict fully global interactions, indicating room for further advancement. Stereo4D  [12] attached a temporal motion head to the DUSt3R architecture used to estimate the position of scene points at any query time between the time of the two input frames and introduced a large dataset of internet videos with noisy point trajectories to train the motion head.
@@ 3.1 Preliminary
Conversely, in a dynamic scene, most point maps cannot be described solely by simple geometric transforms, as they inherently encode scene motion. Note that by leveraging $\prescript{c}{}{\mathbf{X}}^{a}_{t_{1}}$ and $\prescript{c}{}{\mathbf{X}}^{a}_{t_{2}}$, we can effectively analyze the motion of the scene points visible in frame $a$ from time of frame $t_{1}$ to the time of frame $t_{2}$.
@@
In the VGGT method, each input image is first decomposed into image tokens via DINOv2  [23]. These tokens, along with a learnable camera token and four learnable register tokens per image, are then input to a global aggregator module. Since predictions are made in the initial frame’s coordinate system, the camera and register tokens for the first image differ from those for subsequent frames, enabling the network to discriminate among frames. The aggregator module consists of alternating local and global attention layers: local layers facilitate interactions within a single image, while global layers promote interactions across images.
@@ 3.2 Point Tracking and Dynamic Reconstruction
Our method takes a sequence of $N$ frames, $(I_{t})^{N}_{t=1}$, and a query time $q$, and jointly reconstructs the scene and tracks observable points in 3D. To this end, for each frame $I_{t}$, we predict two pixel-wise maps expressed in the coordinate system of the first frame $I_{1}:\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{t}$, capturing point positions at the observation time, and $\prescript{1}{}{\mathbf{\hat{M}}}^{t}_{q}=\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{q}-\prescript{1}{}{\mathbf{\hat{X}}}^{t}_{t}$, capturing the motion of points from time of frame $t$ to the time of query frame $q$. By jointly estimating these maps, our method concurrently reconstructs the 3D structure and tracks point motion to the query frame. Notably, our tracking approach establishes direct correspondences between each frame and the query frame, rather than using pairwise frame-to-frame tracking, enabling efficient capture of overall motion and deformation over extended intervals.
@@ 4.1 Experimental Setups
where $\delta_{3D}=\{\mathrm{0.1m,0.3m,0.5m,1.0m}\}$. Since predictions are made in the coordinate system of the first frame’s camera, we align the estimated 3D positions with the ground truth by scaling them by the median ratio of the norms of the estimated and ground-truth positions. That is, we multiply the predicted coordinates by the $s$ factor:
@@
During training, sequences of 2 to 10 frames are randomly selected from all available sequences. For Kubric Movi-F, we use a stride of $1$; for all other datasets, we use a stride ranging from $1$ to $4$. TartanAir and Virtual KITTI 2 are sampled with half the frequency of the different datasets.
We observe that training with the final tracking objective from the start, using model weights initialized with VGGT’s weights, degrades performance. Thus, we perform training in two phases. In the first phase, we train DePT3R using only the camera, depthmap, and pointmap losses. The intrinsic embedding is used in the first phase, but the query embedding is not. In the second phase, we integrate the query embedding and motion head, along with the tracking loss. To improve training stability and speed, we do not use a confidence-based loss in the second phase.
We employ common data augmentation strategies, such as color jittering, random aspect ratios, and random center crops  [36], applied uniformly to each image in the sequence. We use the Adam optimizer with a warmup phase, followed by cosine scheduling. Gradient checkpointing and bfloat16 precision are utilized to enhance efficiency and reduce GPU memory usage. In the first phase, we train on all datasets, using a learning rate of $5\times 10^{-5}$ for the intrinsic embedding and a learning rate of $5\times 10^{-6}$ for the other weights. In the second phase, we train only on PointOdyssey, DynamicReplica, and Kubric, using a learning rate of $1\times 10^{-5}$ for the motion head and the query embedding, and a learning rate of $1\times 10^{-6}$ for all other weights. Both training phases are run on a single NVIDIA RTX 6000 and take approximately 3 days. We use PyTorch 2.3.1 as our machine learning framework. Experiments are performed on a machine equipped with an AMD Ryzen Threadripper PRO 7975WX 32-core CPU and Ubuntu 22.04 installed.
@@ 4.2 Experimental Results
3D Point Tracking. We first evaluate the 3D point tracking task across three datasets: the test sets of the synthetic datasets PointOdyssey and DynamicReplica, which feature significant camera and scene motion, and PanopticStudio, a real-world dataset with no camera motion. For evaluation, we followed the same setup of St4RTrack [7] and selected 64 consecutive frames from 50 randomly selected sequences in each dataset. To compare with previous work, we restrict our evaluation to the trajectories of the points from the first frame, even though our method can track all points. Following St4RTrack, we exclude PointOdyssey sequences that contain fog or are generated in the Kubric style.
As shown in Table 2, our method significantly outperforms all baselines across all datasets. Remarkably, even though our model was trained only on sequences of at most 10 images, it achieves impressive performance on much longer sequences. Moreover, unlike other approaches that require windowing or global optimization to handle extended sequences due to memory or architectural constraints, our method can generate predictions for a large number of frames in a single forward pass.
3D Reconstruction. Following St4RTrack [7], we diverge from previous work, directly evaluating the accuracy of our estimated 3D reconstruction with the APD and EPE metrics. We use the TUM RGB-D Benchmark, a real-world dataset with significant camera and scene motion, and the PointOdyssey test set. We report metrics on 50 randomly chosen sequences of 64 consecutive frames. For the TUM RGB-D Benchmark, we filter out points with depths between 0.1 and 5 meters, as the depth camera’s accuracy degrades at long ranges. On the TUM RGB-D Benchmark (Table 3), DePT3R achieves an APD of 92.22 and an EPE of 0.0968, competitive with VGGT and significantly better than the other baselines. On PointOdyssey, DePT3R achieves an EPE of 0.0406 and an APD of 98.01, noticeably better than all of the baselines. These results demonstrate that DePT3R achieves state-of-the-art 3D reconstruction accuracy on both synthetic and real-world data, including out-of-distribution scenarios, using synthetic training data.
@@
Figure 4: Qualitative ablation study on the Panoptic Studio dataset. The 3D trajectories of scene points are projected onto the last frame. Without the intrinsic embedding, the model can identify the direction of motion but struggles to accurately place it or gauge its magnitude, leading to significant errors in point trajectory estimation.
@@
Figure 5: GPU memory usage comparison between SpatialTrackerV2, VGGT and our DePT3R method across varying numbers of query points. SpatialTrackerV2 and VGGT exhibit rapid increases in GPU memory consumption, exceeding the 48 GB limit at just 40k and 22.5k query points, respectively. In contrast, DePT3R efficiently handles 268k query points, requiring only 12 GB of memory. All experiments were performed on an RTX A6000 GPU (48 GB). Zoom in for better visualization.
@@ 5 Conclusion and Future Works
Summary. In this paper, we introduce DePT3R, a simple, novel framework for simultaneously performing dense 3D reconstruction and point tracking of dynamic scenes from unposed monocular video. Despite training exclusively on synthetic datasets, our approach achieves strong performance on challenging real-world benchmarks, highlighting the inherent synergy between reconstruction and point-tracking tasks. Unlike prior methods, DePT3R leverages global temporal attention to enhance both accuracy and computational efficiency, enabling dense tracking of all visible points in a single forward pass. Additionally, we effectively leverage intrinsic information to significantly improve reconstruction and motion estimation.
@@ Acknowledgements
This work was supported by Professor M. Khalid Jawed and, in part, by the US Department of Agriculture (Grant numbers 2024-67021-42528 and 2022-67022-37021).